OpenStudy (anonymous):

is 0.315 with line over 315 a irratiional number or rational

OpenStudy (saifoo.khan):

o.O

myininaya (myininaya):

rational

OpenStudy (anonymous):

RATIONAL BECAUSE YOU CAN DERIVE A FRACTION OUT OF REPEATING DECIMALS.

myininaya (myininaya):

let x=.315315315.... then 1000x=315.315315315.... 1000x-x=315 999x=315 x=315/999 this is a rational number since we can write it an integer over an intger

OpenStudy (zarkon):

\[.\overline{315}=\frac{315}{999}\]